one half of louisa's age three years from now increased by three-quarters of her age three years agois 8. How old is louisa now?
Answer by stanbon(75887)

one half of louisa's age three years from now increased by three-quarters of her age three years agois 8. How old is louisa now?
---
Equation:
(1/2)(L+3)+(3/4)(L-3) = 8
----
Multiply thru by 4 to get:
2(L+3) + 3(L-3) = 32
5L - 3 = 32
5L = 35
Louisa's age now = 7 yrs
================================
Cheers,
Stan H.